• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name K. Grimm

    Log entry time 16:02:28 on October19,2009

    Entry number 295717

    keyword=Day Shift Summary

    SL: Klaus Grimm
    TO: Luis Mercado
    3rd:Zafar Ahmed


    Production runs at 97uA
    13397 44.9 IHWP IN, reached 1028k good events
    13398 43.6 IHWP OUT
    13399 42.8
    13400 42.1
    13402 40.5
    13404 43.7
    13405 46.2

    Pedestal runs (no beam)


    Shift starts with Happex run 13397 at 97uA and IHWP IN

    08:21 Compton GSO HV trips. Turned HV on (was set to -1500V from previous shift)
    08:41 Changed IHWP to OUT. Starting slug 27
    08:43 Starting Happex run 13398
    08:43 Starting Compton run 20513
    08:56 MCC is going to try to adjust pathlength to see if that can improve the
    Compton photon background rate.
    09:24 Spot++ run 25082 on left arm
    Raster size on target: 6 x 5.5 mm
    09:40 Start Happex run 13399
    09:56 Stopped Compton run 20513 per request of Alex. He is at MCC taking over the Compton DAQ for pathlength tuning.
    10:02 Compton GSO HV trips. Turned HV on. Somebody changed during this morning the default HV remotely from -1500V back
    to -1800V. OK, Alex is remotely controlling/changing Compton GSO HV settings at MCC.
    10:38 Start Happex run 13400
    10:50 Compton DAQ has been handed over to shift crew
    10:54 Start Compton run 20516
    11:34 Happex DAQ crashed, run 13401 is junk
    11:38 Start Happex run 13402
    12:35 Start Happex run 13403 (mostly no beam, good for pedestal check)
    12:37 no beam. MCC: the east arc BPMs went into alarm simultaneously causing a fast shutdown that would not reset.
    12:48 Compton reboot by Alex
    12:49 Start Compton run 20517 (w/o beam)
    13:25 MCC is ready to send beam
    13:26 Start Compton run 20518
    13:26 Start Happex run 13404
    13:37 Spot++ run 25089, raster size on target seems unchanged.
    Kai has taken over the left arm DAQ for testing alayser code
    13:59 Compton GSO HV trip. Turned HV on (-1800V)
    14:01 Compton GSO HV trip. Turned HV on (-1800V)
    14:26 Start Happex run 13405
    15:20 beam is off due to dogleg 2 trip
    15:23 Start Happex run 13406 (~20min no beam)